# PVPO: Pre-Estimated Value-Based Policy Optimization for Agentic Reasoning

### v1

@@ 6.3.1 Convergence Speed
6.3.2 Total Time Consumption
[…]
Figure 4: Training stability study of PVPO on multi-step retrieval datasets.
6.4 Stability Evaluation
We tracked PVPO training metrics to show its stability. Figure 4 (a) shows PVPO achieves a much higher average reward than GRPO. With similar KL divergence in Figure 4 (b), this improvement comes not from more aggressive updates, but from better gradient direction estimates. As shown in Figure 4 (c), PVPO has much lower advantage variance, leading to more reliable and consistent update directions. PVPO also maintains exploration without losing stability. Figure 4 (d) shows it keeps higher policy entropy under a similar KL constraint, which helps avoid premature convergence to a local optimum. Overall, PVPO addresses key problems in RL: it supports high exploration, low variance, and high rewards. This evidence answers Q4: PVPO achieves more stable training than existing methods.
6.5 Case Study: Low Sampling Budget
We ran a case study to address Q5. We reduced the number of rollouts from 5 (used in the main experiments) to 2. For comparison, we report GRPO’s performance with a full budget. Figure 5 (a) shows that PVPO with a low budget remains close to the fully budgeted GRPO. We calculate computational cost by multiplying the number of rollouts with the average number of tool calls in trajectories. As shown in Figure 5 (b), PVPO’s average cost is only 4.3, which is much lower than GRPO’s 11.7. PVPO achieves 97% of GRPO’s performance (55.0% vs. 56.8%) while using less than 40% of the computational cost. This strong sample efficiency comes from the high-quality, low-variance training signals provided by Static V Estimate. The model can update its policy efficiently using fewer rollouts. This answers Q5: model training clearly benefits from Static V Estimate.
Figure 5: Low sampling budget study of PVPO on multi-step retrieval datasets. The NN denotes the number of trajectories in each single rollout. NN=5 is the full budget and NN=2 is the low budget.
7 Conclusions
In this paper, we propose PVPO, an efficient critic-free reinforcement learning algorithm designed to optimize policy learning for complex tasks. By introducing a Static V Estimate as an external advantage reference and integrating it with group sampling for effective data filtering, PVPO addresses the limitations of extensive sampling and biased intra-group comparisons inherent in prior methods. Our approach yields stable, low-variance training signals, accelerates convergence, and significantly reduces computational costs. Extensive experiments across nine diverse benchmarks in multi-hop question answering and mathematical reasoning demonstrate that PVPO achieves state-of-the-art performance and strong generalization, even with small-scale models and limited resources. PVPO introduces substantial improvements in reasoning and tool use, supports scalable training, and ensures consistent performance, thereby demonstrating strong potential for widespread real-world application.

### v3

@@ A.2 Group Sampling Analysis
A.3 Additional Experiment Results
To further verify the scalability of our proposed PVPO method, we conduct integration experiments on multi-hop QA tasks. Specifically, we combine PVPO with the sequence-level importance ratio module proposed in GSPO and remove the KL loss constraint as introduced in DAPO. The results, shown in Table 4, demonstrate that PVPO not only provides strong baseline improvements over GRPO, but also achieves further performance gains when integrated with these advanced RL methods. In particular, the combination with DAPO (w/o KL) yields the best accuracy and LasJ scores, while integration with GSPO’s sequence-level importance ratio also presents consistent improvements. In particular, the combination with DAPO (w/o KL) yields the best accuracy and LasJ scores, but also incurs significantly more tool calls (8.14 per query), resulting in greater inference costs. By contrast, GSPO’s sequence-level importance ratio offers improvements with relatively lower tool call overhead (2.19 per query). Therefore, the trade-off between performance and inference cost should be considered when choosing an integration strategy for different practical scenarios. These findings confirm that PVPO is highly compatible and complementary when used alongside other state-of-the-art RL algorithms.
Table 4: Experimental results of PVPO’s orthogonal integration with SOTA RL algorithms (DAPO, GSPO) and scalability evaluation on multi-hop QA tasks. “w/ Seq-Ratio” refers to the sequence-level importance ratio from GSPO, and “w/o KL” means removing the KL loss constraint as in DAPO.
Method Average
Acc LasJ ToolCalls
GRPO-ReSearch 48.6 58.0 2.46
PVPO-ReSearch 54.4 62.6 2.96
w/ Seq-Ratio (GSPO) 55.1 62.4 2.19
w/o KL (DAPO) 58.8 67.1 8.14